OpenStudy (anonymous):

100 plus the product of a number and -2 equals 50. What is the number?

OpenStudy (anonymous):

100+(-2)*x=50 100-2x=50 100-50=2x 50=2x 25=x

OpenStudy (anonymous):

thank you